OpenStudy (anonymous):

sum of (n^10 +10)/n! n=1 to infi

OpenStudy (anonymous):

i tired solving it as a(n+1) = (n+1)^10 +10 /(n+1)! and an = n^10 + 10/n!

OpenStudy (anonymous):